OpenStudy (anonymous):

35sinx-100cosx+150cos2x=0 , i need to get values of x.

OpenStudy (michele_laino):

sorry is it \[\cos(2x)\] or \[(\cos x)^{2}\]?

OpenStudy (anonymous):

cos(2x)

OpenStudy (michele_laino):

I rewrite your equation as below: \[7 \sin x-20 \cos x +30 (cosx)^{2}-30 (\sin x)^{2}=0\] from which: \[\cos x[(7 \tan x-20)+30 \cos x(1-(\tan x)^{2}]=0\]

OpenStudy (jhannybean):

Oh you divided everything by 5 first.

OpenStudy (michele_laino):

now, applying the canceling law of product, we have: cos x=0 or\[(7 \tan x-20)+30 \cos x (1-(\tan x)^{2}))=0\]

OpenStudy (michele_laino):

@Jhannybean that's right!

OpenStudy (michele_laino):

from cos x=0, we have our first solutions: namely: \[x=\frac{ \pi }{ 2 }+k \pi\] where \[k=0, \pm1, \pm2,\pm3....\]

OpenStudy (michele_laino):

after that, I don't know!

OpenStudy (anonymous):

In This Part 30cosx[1-(tanx)^2] shouldnt u take [cosx]^2 as common factor not cosx

OpenStudy (jhannybean):

He did, but he factored it out of the whole equation rather than just that portion.

OpenStudy (michele_laino):

we can try to use the subsequent identity: \[\frac{ 1 }{ \cos x }=\sqrt{(\tan x)^{2}+1}\]

OpenStudy (michele_laino):

so our second equation, can be rewritten as below: \[7 \tan x-20+30\frac{ 1-(\tan x)^{2} }{ \sqrt{1+(\tan x)^{2}} }=0\] and solving for tan x

OpenStudy (anonymous):

Im Inclined to take the first equation answer , but what would be the value for tanx ?

OpenStudy (michele_laino):

@AmrAhmed please set tan x=z in my second equation, and you will get an irrational equation in z

OpenStudy (michele_laino):

here is your equation in z: \[(7z-20)\sqrt{1+z ^{2}}=30z ^{2}-30\]

OpenStudy (anonymous):

Well it still doesn't solve it since it gives me areas irregularities in the diagram since x is an intersection point ,good job however .

OpenStudy (michele_laino):

after squared both sides of the equation above I will get: \[851 z ^{4}+280 z ^{3}-2249 z ^{2}+280 z +500=0\] which can be soved using tangents method of Newton, for example

OpenStudy (michele_laino):

thank you! for your appreciation @AmrAhmed

OpenStudy (michele_laino):

thank you too @Jhannybean

OpenStudy (jhannybean):

Np:)